OpenStudy (anonymous):

For the reaction given below, how many liters of ammonia form when 7.5 L of hydrogen reacts? equation: N2(g)+ 3 H2(g) yields 2 NH3(g)

OpenStudy (anonymous):

This is assuming the question is talking about at STP. But for every mole of a substance there is 22.4 liters of it at STP. So the numbers before the compounds in the equation are mole ratios. For every 3 moles of H2 there are 2 moles of NH3. We know we have a total of 7.5L of H2. So we could mathematically set it up like this: \[\frac{ 7.5L H2 }{ 1 } \times \frac{ 2 moles NH3}{3 moles H2} =\frac{ 7.5 \times 2 }{ 1 \times 3 } =\frac{15}{3}=5 L NH3\]

OpenStudy (anonymous):

thx man